# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v1

@@ A.1 List of Parameters
$\beta=0.985$
Discount factor
$\gamma=1.45$
Elasticity of marginal utility
$\alpha=0.3$
Output elasticity of capital
$\delta=0.1$
Annual depreciation rate of capital
@@
$\zeta=0.0021$
Contribution rate of carbon emissions to temperature
$\sigma_{i,t}$
Carbon intensity (GtC per trillion \$)
$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$
Abatement cost parameters
$\pi_{1,i}$, $\pi_{2,i}$
Climate damage parameters
@@
$g_{i,0},g_{i,t}$
Growth rate of TFP
$d_{i}$
Parameters of TFP growth
@@ A.2 List of Countries
Africa
Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
European Union [^13] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience.
Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
Eurasia
Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
Latin America
Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St. Kitts and Nevis, St. Lucia, St. Vincent and the Grenadines, Suriname, Trinidad and Tobago, Uruguay, Bolivarian Republic of Venezuela, and the U.S. Virgin Islands.
@@
Middle East
Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
Other Non-OECD Asia
Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
Other High-Income
Australia, Canada, Republic of Korea, New Zealand, Singapore.

### v2

@@ A.1 List of Parameters
$\beta=0.985$ Discount factor
$\gamma=1.45$ Elasticity of marginal utility
$\alpha=0.3$ Output elasticity of capital
$\delta=0.1$ Annual depreciation rate of capital
@@
$\zeta=0.0021$ Contribution rate of carbon emissions to temperature
$\sigma_{i,t}$ Carbon intensity (GtC per trillion \$)
$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$ Abatement cost parameters
$\pi_{1,i}$, $\pi_{2,i}$ Climate damage parameters
@@
$g_{i,0},g_{i,t}$ Growth rate of TFP
$d_{i}$ Parameters of TFP growth
@@ A.2 List of Countries
Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
European Union [^17] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience. Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
Latin America Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St. Kitts and Nevis, St. Lucia, St. Vincent and the Grenadines, Suriname, Trinidad and Tobago, Uruguay, Bolivarian Republic of Venezuela, and the U.S. Virgin Islands.
@@
Middle East Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
Other High-Income Australia, Canada, Republic of Korea, New Zealand, Singapore.
@@ A.8.5 Regional MAC and SCC under Alternative Emission Caps
A.8.6 Sensitivity Analysis on Marginal Abatement Cost
In Figure A.15, we provide additional sensitivity analysis results on the abatement cost. Recall that, the MAC in the model is
$\mathrm{MAC}{}_{i,t}=1,000b_{2,i}\mu_{i,t}^{b_{2,i}-1}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t)).$
We investigate a scenario wherein the MACs across all regions experience a variation of ±20% from the baseline MAC levels at the same emission control rates $\mu_{i,t}$. Specifically, we explore the effects of adjusting two key parameters $b_{1,i}$ and $b_{3,i}$, by ±20% from their baseline values. The result shows that the equilibrium permit price is closely linked to the level of the MACs. When the MACs are elevated across all regions, the permit price also experiences a corresponding increase with almost identical changes of the MACs, and vise versa. For example, in 2040, the baseline MAC in the US is \$556, it is changed to \$644 or \$467, and the permit price is changed from \$404 to \$492 or \$315. That is, both the MACs and the permit price increase \$88 or decreases \$89 in 2040. At the same time, the regional SCC has little changes, as the existence of the emission caps makes the future temperature have little changes and the changes in the MACs have little impact on marginal climate damages in future periods and social discount rates. This again verifies our relation between the optimal carbon tax, the MAC, and the permit price (i.e., equation (20)).
Figure A.15: Sensitivity Analysis on the MAC.